Mary Inc. is considering mutually exclusive Projects A and B, whose cash flows are shown below. If the decision is made by choosing the project with the higher IRR, will there be any value loss due to the IRR-based decision? If there is a loss, how much value will be forgone? The WACC is assumed to be 9.5%.

WACC:

Year

9.5%

0

1

2

3

4

CFA

−$2,020

$730

$730

$740

$740

CFB

−$4,100

$1,400

$1,500

$1,520

$1,530
Answer #1

1 Project A TIME Year 0 Year 1Year 2Year 3 Year 4 3 IRR (in %) 4 Annual Cash Flows (ACFs) 5 PV Factor@ IRR 6 PV of ACFs 7 NPV 16.85 730 0.856 2020.00624.72 740 0.627 534.62 463.78 730 740 0.536 396.89 2020 1.000 0.732 0.00 9 Project B 10 11 IRR (in %) 12 Annual Cash Flows (ACFs) 13 PV Factor IRR 14 PV of ACFs 15 NPV TIME Year 0 Year 1Year 2 Year 3 Year 4 16.49 -4100 1530 0.543 -4100.00 1201.86 1105.47 961.67 831.00 1400 1500 1520 1.000 0.858 0.737 0.633 0.00

Using EXCEL's Goal Seek Function to solve for IRR (the WACC at which NPV is zero), we observe that Project A's IRR is greater (at 16.85 % ) and hence the same should be selected.

1 Project A TIME Year 0 Year 1Year 2 Year 3 Year 4 3 WACC (in %) 4 Annual Cash Flows (ACFs) 5 PV FactorIRR 6 PV of ACFs 7 NPV 9.50 2020 1.000 2020.00 730 0.913 666.67 730 0.834 608.83 740 0.762 563.62 740 0.696 514.72 333.84 9 Project B 10 11 WACC (in %) 12 Annual Cash Flows (ACFs) 13 PV Factor IRR 14 PV of ACFs 15 NPV TIME Year 0 Year 1 Year 2 Year 3 Year 4 9.50 1530 0.696 -4100.00 1278.54 1251.02 1157.71 1064.23 -4100 1400 1500 1520 1.000 0.913 0.834 0.762 651.50

1 Project A TIME Year 0 Year 1 Year 2 Year 3 Year 4 WACC (in %) 4Annual Cash Flows (ACFs) 5 PV Factor IRR 6 PV of ACFs 7 NPV 8 9 Project B 10 11 WACC (in %) 12 Annual Cash Flows (ACFs) 13 PV Factor@ IRR 14 PV of ACFs 15 NPV 9.5 2020 730 730 740 740 (C4 C5) (G4G5) -SUM(C6:G6) TIME Year 0 Year 1 Year 2 Year 3 Year 4 9.5 4100 1400 (1/(1+(SB$11/100 =(D12*D13) 1500 1520 1530 (F13/(1+(SB$11/100) G12 G13) (D13/(1HSB$11/100)1 (E13/1+(SB$11/100 (C12*C13) (E12 E13) =(F12*F13) -SUM(C14:G14)

As is observable NPV (B) > NPV(A). This implies a loss in value if the IRR criterion is used to select the project, as Project A would be selected in such a case even though it has a lower NPV.

Value Loss = NPV(B) - NPV(A) = 651.5 - 333.84 = $ 317.65